1. 2ln 12-ln 9
2. 1/4ln3+1/4lnx
3. ln(x+2/3)=12
4. 4e^3x+1.2=14
5.Protactinium-234m, a toxic radioactive metal with no known use has a half-life of 1.17 minutes. How long does it take for a 10mg sample to decay to 2mg?

Answer by stanbon(75887)

1. 2ln 12-ln 9 = ln[12^2/9] = ln[16] = 2.7726
----------------------------------
2. (1/4)ln3 + (1/4)lnx = ln[3^(1/4)*x^(1/4)] = ln[(3x)^(1/4)]
-------------------
3. ln(x+2/3)=12
Is this what you mean??
x + 2/3 = e^12
Or, is this what you mean??
(x+2)/3 = e^12
========================
4. 4e^3x+1.2=14
Again, it is not clear what you mean.
Use parentheses.
-----------------------------
5.Protactinium-234m, a toxic radioactive metal with no known use has a half-life of 1.17 minutes. How long does it take for a 10mg sample to decay to 2mg?
--------
2 = 10(1/2)^(x/1.17)
----
(1/2)^(x/1.17) = 1/5
----
Take the log::
(x/1.17)log(1/2) = log(1/)5
x/1.17 = log(1/5)/log(1/2)
x/1.17 = 2.3229
x = 2.72 minutes
----------------------
Cheers,
Stan H.
----------------